OpenStudy (anonymous):

integral of sin(1/x)dx between 0 to 1 converges, why?

OpenStudy (anonymous):

(Improper integral of course) :)

OpenStudy (turingtest):

let u=1/x

OpenStudy (anonymous):

It does not work with u substitution.. on WolframAlpha it's something which I have never seen :(

OpenStudy (turingtest):

oh yeah, sorry, just realized I messed up :P

OpenStudy (anonymous):

The integral from 0 to 1 DOES converge, but I cant understand why.. (any comparasion test or something..?)

Parth (parthkohli):

An integral is a sum. A sum converges when?

Parth (parthkohli):

Can only provide an intuitive picture for now.

OpenStudy (anonymous):

But a sum works only for Natural numbers.. and where n (the index) starts from? The integral is from 0 to 1, so the sum is from 0 to infinity? [wolfram says the sum diverges. I don't know if in this case you can use sums :X

Parth (parthkohli):

Again, I can only provide a picture to you. sin will never exceed 1, that I am sure of.

Parth (parthkohli):

But this is an interesting question. Let me think about it.

OpenStudy (turingtest):

Not a great proof of anything, but as a comparative insight, you can notice that\[-1\le\sin\frac1x\le1, x\in\mathcal R\]so the function will never go off to infinity, as I'm sure you already realized. Pictorially the function starts oscillating like crazy as \(x\to0\), but there will always be a curve below the x-axis to cancel the one above in that region, so the integral should be tend to zero around x=0. As you go out towards x=1, you are clearly defined and bounded, so you get some finite answer.

OpenStudy (anonymous):

Intuitively you are 100% right, but is there any mathematical method to assure this? Sometime intuition in math does not work :)

OpenStudy (turingtest):

I'm trying to conjure up something more concrete, but so far all I have is that...

Parth (parthkohli):

Related: 1 - 1 + 1 - 1 ... = 0.5 Some people argue that this has no definite value while some agree that it is 0.5.

Parth (parthkohli):

The problem here is that there is no limit to sine as the input approaches infinite. http://math.stackexchange.com

OpenStudy (anonymous):

@oshrinuri , the limits are flawed. Aren't they?

OpenStudy (anonymous):

No, it's from 0 to 1. They didn't ask us what is the value of the integral, only whether it converges / diverges..

OpenStudy (anonymous):

\[\large \int_0^1 \frac{\sin x}{x}dx= \lim_{R \to 0} \int_R^1 \frac{\sin x}{x}dx= \lim_{R \to 0 } \left. -\frac{\cos x}{x} \right|^1_R - \int_R^1 \frac{\cos x}{x^2}dx \]

OpenStudy (anonymous):

\[ \lim_{x \to 0 } \frac{\cos x}{x}=\frac{1}{x \to 0}\sim \infty \]

OpenStudy (turingtest):

but it's\[\int_0^1\sin\frac1xdx\]

OpenStudy (anonymous):

oh lord, that makes sense then ^^ thank you.

OpenStudy (anonymous):

stupid question :'(

OpenStudy (turingtest):

can we just say\[-\int_0^1dx\le\int_0^1\sin\frac1xdx\le\int_0^1dx\]? XD seems silly, but this would suggest that it converges, right?

OpenStudy (anonymous):

another way would be the taylor expansion, studying the radius of convergences and see that it lies within the given bounds of integration.

OpenStudy (anonymous):

nevertheless, the above approximation holds true, in fact you can use absolute convergence: \[\large 0\leq\left| \sin \left( \frac{1}{x} \right) \right| \leq 1 \] and use that absolute convergences implies convergences. Which is what @TuringTest did, holds for approximations for sums in special

OpenStudy (anonymous):

Thanks guys! :D